# Warm-Starting Collision-Free Model Predictive Control With Object-Centric Diffusion

### v1

@@ I INTRODUCTION
Generating collision free and dynamically feasible motions remains a central challenge in robotic manipulation, especially in cluttered and non convex environments. Classical sampling based planners explore high dimensional spaces effectively but do not handle dynamics [lavalle_rapidly-exploring_1998, karaman_sampling-based_2011], while trajectory optimization and MPC based methods explicitly incorporate smoothness, dynamics and collision constraints [gaertner_collision-free_2021, haffemayer_collision_2025]. Despite recent progress in collision free MPC, these solvers remain highly sensitive to initialization and often converge to local minima in complex scenes.
Learning-based trajectory priors have recently emerged as a way to address this limitation. In particular, diffusion models [ho_denoising_2020] have shown strong ability to capture the multimodal structure of feasible trajectories, where several distinct solutions may exist around obstacles. When combined with trajectory optimization, these priors help guide the solver toward promising regions of the search space [seo_presto_2025, huang_diffusion-based_2023]. However, to generalize to unseen environments, diffusion models must be conditioned on scene information [huang_diffusionseeder_2024], and designing an effective and compact representation of obstacles remains a key challenge.
@@
From this perspective, we first argue that an object-centric representation of the scene provides a compact and highly informative conditioning signal for trajectory diffusion. We propose a Slot-Attention–based object-centric encoding that extracts a small set of latent obstacle descriptors from estimated object poses [locatello_object-centric_2020]. This representation captures the essential collision structure while enabling fast inference and improved generalization compared to alternative conditioning strategies. We then show that the resulting scene-conditioned diffusion model interacts naturally with a constraint-based MPC solver, which rigorously and efficiently projects the diffused trajectory prior onto a dynamically feasible and collision-free trajectory. In contrast to previous approaches that rely on simple gradient-based post-processing of diffusion samples, our diffusion prior is refined through full OCP projection running in real time on a physical robot. In addition, we introduce a systematic dataset-generation procedure that combines sampling-based planning with constrained trajectory optimization to produce high-quality supervision trajectories.
@@ II-A Classical and learning-based motion planning methods
Classical sampling-based planners [reif_complexity_1979, canny_complexity_1988], explore high-dimensional configuration spaces effectively but do not account for system dynamics [lavalle_rapidly-exploring_1998, lamiraux_prehensile_2022]. Trajectory optimization methods [ratliff_chomp_2009, kalakrishnan_stomp_2011, schulman_motion_2014] address this limitation by explicitly incorporating smoothness and dynamic constraints. MPC further extends this approach, formulating collision avoidance as either a hard constraint or a penalization in the optimal control problem [mayne_model_2014, gaertner_collision-free_2021, haffemayer_collision_2025], allowing for online planning of dynamically feasible, collision-free motions. Alternative approaches such as Model Predictive Path Integral control [sundaralingam_curobo_2023, legged_xue_2025] and tree-search methods like Monte Carlo Tree Search [bonanni_monte_2025] introduce stochastic exploration to mitigate local minima, but their computational cost often limits real-time applicability. Despite these advances, classical TO, MPC, and sampling-based extensions remain sensitive to initialization and prone to local minima in complex, non-convex environments.
Learning-based methods provide a complementary solution by leveraging prior experience from demonstrations [kober_imitation_2010, mandlekar_what_2021] or previously successful trajectories [ichter_learning_2019, qureshi_motion_2019, urain_learning_2022, jordana_infinite-horizon_2025], guiding planners or optimizers toward promising regions of the search space. In particular, providing a high-quality warm-start near a feasible solution can dramatically accelerate convergence and improve final solution quality [stolle_policies_2006, mansard_using_2018, lembono_learning_2020]. Achieving such warm-starts is challenging due to the high degree of multi-modality in robotic trajectory generation, where multiple distinct solutions may exist for the same task.
@@ II-B Diffusion models in robotics
Diffusion models [ho_denoising_2020] have recently emerged as a powerful tool to capture the multi-modal distributions of feasible trajectories. Early work such as Diffuser [janner_planning_2022] iteratively denoises trajectories using a learned model and differentiable guiding functions, while Diffusion Motion Planning [carvalho_motion_2023] adds gradient-based guidance for task-specific objectives, such as collision avoidance or smoothness. Extensions incorporating model-based projections [romer_diffusion_2025] further improve the quality of generated trajectories. Composable Diffusion Models [power_sampling_2023] learn distributions over constraint-satisfying trajectories and leverage them to handle novel combinations of constraints. While effective in low-dimensional or structured problems, these approaches generally struggle to generalize to cluttered or unstructured scenes. [huang_toward_2024] recently applied diffusion models to approximate globally optimal nonlinear MPC. Their approach generates a diverse offline dataset of optimal control sequences, trains a diffusion model to capture the multi-modal distribution of solutions, and samples near-globally optimal trajectories at runtime without relying on initial guesses. While effective in simplified simulated environments, it has not been demonstrated on real robots.
@@ II-C Scene-conditioned diffusion approaches
To improve generalization to unseen environments, several works condition diffusion models directly on scene information [chisari_learning_2024, tian_warm-starting_2025, huang_diffusion-based_2023, seo_presto_2025]. Diffusion Seeder [huang_diffusionseeder_2024] goes a step further by deploying diffusion-generated trajectories on a real robot; however, like prior work such as [carvalho_motion_2023], it generates only open-loop trajectories that are tracked by a low-level controller. These approaches do not perform closed-loop replanning, enforce dynamic feasibility during generation, or integrate with a receding horizon MPC framework. While they improve robustness compared to purely simulated methods, they remain limited to simplified or low-dimensional setups and do not demonstrate real-time, collision-aware control with compliance.
@@ II-D Environment representation
A key factor in scene-conditioned planning is how the environment is represented. Image-based representations [chi_diffusion_2024] can lose geometric fidelity, whereas point clouds [chisari_learning_2024, tian_warm-starting_2025] capture 3D geometry but are high-dimensional, unstructured, and require specialized networks to extract object-level features. Similarly, Diffusion Seeder [huang_diffusionseeder_2024] relies on depth images, which provide direct 3D information but also suffer from high dimensionality and noise, limiting inference speed and generalization. Structured scene representations, such as PRESTO [seo_presto_2025], encode key robot configurations labeled for collision status, while object-centric approaches like Slot-Attention [locatello_object-centric_2020] map inputs into distinct object slots. These structured representations yield compact latent spaces, better generalization, and faster inference, especially in cluttered or unseen environments [chapin_object-centric_2025]. In this work, we adopt a Slot-Attention-based approach and systematically compare it with image-based, PRESTO, and occupancy grid representations (Sec. IV-C). Building on these principles, our approach integrates object-centric, scene-conditioned diffusion priors with a full-dynamics MPC framework. Unlike prior work, we demonstrate real-time, closed-loop control of a physical robot in cluttered, dynamic environments while strictly enforcing collision and dynamic constraints, highlighting the practical benefits of combining generative priors with model-based control.
@@ III DIFFUSION-GUIDED MPC
Figure 3: Object-centric scene representation using Slot Attention. Given an RGB image of the scene, a pre-trained Slot-Attention encoder [locatello_object-centric_2020] extracts a set of object-centric latent embeddings, called slots, each corresponding to a distinct object or obstacle in the environment. These slots are then used to construct a structured representation of the scene, which serves as input to the diffusion-based trajectory generator. This conditioning enables the model to generate trajectories that account for obstacle geometry and layout, while maintaining generalization to unseen environments.
@@ III-A1 Trajectory diffusion
To generate high-quality warm-start trajectories, we train a conditional diffusion model as a learned prior over feasible motion sequences [carvalho_motion_2023, seo_presto_2025]. Our model builds on the Denoising Diffusion Probabilistic Model (DDPM) framework [ho_denoising_2020], which captures complex, multi-modal distributions through a forward noising and reverse denoising process. A robot trajectory $\bm{Q}\!=\!(\bm{q}_{0},\dots,\bm{q}_{T})$ is treated as a sample from an expert distribution. A forward (noising) process gradually adds Gaussian noise to the trajectory over $K$ steps according to a fixed variance schedule $\{\beta_{i}\}_{i=1}^{K}$. The noisy trajectory at step $i$ is given by:
@@ III-A2 Model Architecture
We use a Diffusion Transformer (DiT) that denoises a trajectory $\bm{\hat{Q}}_{i}$ conditioned on the task context. The context includes the robot’s initial configuration $\bm{q}_{0}$, the Cartesian goal $\bm{p_{\text{goal}}}$, and a scene encoding. The latter is provided by a Slot Attention encoder [locatello_object-centric_2020], which converts an RGB image of the workspace into object-centric latent embeddings (slots). This representation captures obstacle information in a compact form, enabling the DiT to generate trajectory proposals that are tailored to each problem instance. We describe the encoder in more detail below.
@@ III-A3 Slot Attention-Based Scene Encoder
The Slot Attention encoder [locatello_object-centric_2020] extracts object-centric latent representations from single-view RGB images of the robot workspace. This structured decomposition of the scene into obstacles provides a compact conditioning signal for the diffusion model, enabling generalization across cluttered layouts without relying on fixed spatial encodings or manual annotation.
@@ III-B2 MPC costs and constraints
d. Collision constraints. To guarantee safety, we impose a hard collision-avoidance constraint at each timestep following [haffemayer_model_2024], requiring the signed-distance $d(\cdot,\cdot)$ between any pair of potentially colliding bodies $\mathcal{B}$ to be greater than a safety margin $\epsilon_{\text{safe}}$:
@@ III-B3 Warm-starting the solver
The performance of the generative model is critically dependent on the quality of its training data. To learn a useful motion prior, the model must be trained on a diverse dataset of expert trajectories that are not only collision-free but also dynamically smooth and efficient. Trajectories from purely geometric planners like RRT [lavalle_rapidly-exploring_1998] are often jerky and not directly executable. Therefore, a dedicated data-generation pipeline is required to produce high-quality examples. We construct our training dataset by generating thousands of random scenes in simulated environments described in Sec. IV. For each scene, we have computed multiple expert trajectories. We first sample a start/goal pair and generate a geometrically valid path using RRT. This path is then used to initialize and solve the OCP defined in (3), resulting in a smooth, feasible expert trajectory. In parallel, the Slot Attention algorithm was trained beforehand on a task-specific dataset. The diffusion model is then trained on this dataset by minimizing the loss in (2).
Figure 6: Impact of the cost gradient-based guidance. Comparing the impact of the cost gradient-based guidance at each diffusion step, as used in [carvalho_motion_2023]. Results show that guidance is not helpful with a low computational budget, given by control requirements.
(a) Success rate
(b) Computation time (ms)
@@ IV-A2 Metrics and protocol
We evaluate our approach against a range of classical and learning-based motion-planning methods to demonstrate that combining diffusion-based warm-starting with OCP refinement yields superior computational efficiency and higher-quality trajectories. The first learning-based baseline is PRESTO [seo_presto_2025], a diffusion model operating in configuration space, where the scene is represented by sampling a predefined set of robot configurations. For each configuration, a binary collision indicator is computed, and the resulting indicator vector is concatenated to form the conditioning input. The second baseline is Motion-Planning Diffusion (MPD) [carvalho_motion_2023], a learning-based approach where an unconditional diffusion model learns to generate motion trajectories, with collision-avoidance encouraged through a cost gradient-based guidance within the diffusion process. For fairness, we adapt both of their architecture and training procedure to match our dataset and task setup. We then compare our approach to both classical and optimization-based planners. RRT Connect [rrt-connect_kuffner_2000] is a widely-used bidirectional sampling-based planner, implemented in Python with C++ bindings, that incrementally builds trees from the start and goal configurations and attempts to connect them. CuRobo [sundaralingam_curobo_2023] is a recent GPU-accelerated framework that combines trajectory optimization with sampling strategies to efficiently handle high-dimensional planning problems. To ensure a fair comparison under the same time budget, we terminate all algorithms after a fixed number of iterations.
Finally, to demonstrate the benefit of warm-starting the OCP, we compare our method to a standard OCP solver [mastalli_crocoddyl_2020], which solves the full optimal control problem defined in (3) using Differential Dynamic Programming (DDP) [leondes_differential_1973, tassa_synthesis_2012] without any warm-start. This method directly optimizes dynamic feasibility and task objectives but remains highly sensitive to local minima.
@@ IV-C2 Impact of the cost gradient-based guidance
Fig. 6 evaluates the impact of adding a cost gradient-based guidance toward the feasible manifold at each denoising step, following [carvalho_motion_2023]. Without OCP refinement, the cost gradient-based guidance improves success at low budgets by steering trajectories closer to feasibility early on. However, when OCP refinement is applied, its benefit diminishes and even harms performance at higher budgets due to unnecessary overhead. We thus omit the cost gradient-based guidance in our final pipeline.
@@ IV-C3 Trade-off between computation and warm-start quality
Experiments were conducted on a Franka Emika Panda robotic arm controlled via torque commands. The MPC framework was implemented in Python with C++ bindings and executed on an AMD Ryzen 9 5950X CPU at $3.4\text{\,}\mathrm{GHz}$. The dynamics of the system were computed using pinocchio [carpentier_pinocchio_2019], and in our setup the dynamics are time-invariant. The optimal control problem (OCP) was solved using mim-solvers [jordana_stagewise_2025] inside crocoddyl [mastalli_crocoddyl_2020], consistent with the simulation setup described in Section IV-A. The robot dynamics were modeled using inertial parameters from [gaz_dynamic_2019]. Obstacle detection and pose estimation were performed in real-time using AprilTags [olson_apriltag_2011], with known obstacle dimensions and external-camera-based object tracking. Control commands were computed at $1\text{\,}\mathrm{kHz}$, while the OCP was solved at $100\text{\,}\mathrm{Hz}$ with a prediction horizon of $N_{h}=15$ nodes and a time step of ${\Delta t}=$10\text{\,}\mathrm{ms}$\$, resulting in an optimization horizon of $150\text{\,}\mathrm{ms}$. At each OCP iteration, we compute the optimal trajectory and the corresponding Riccati gains [dantec_first_2022]. The Riccati gains are used to apply a low-level feedback policy at $1\text{\,}\mathrm{kHz}$ until the next replanning time. A collision safety margin of $1\text{\,}\mathrm{cm}$ was enforced via a lower-bound constraint on the signed-distance. The MPC operated under a receding-horizon scheme. Whenever a new high-level task was issued (e.g., reaching a new target position), the diffusion model generated a warm-start trajectory, replacing the previous solution. As Fig. 8 shows, this warm-started OCP enabled reactive and efficient collision-avoidance in cluttered and dynamically changing environments.

### v2

@@ Abstract
publicationid: pubid: © 2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I INTRODUCTION
Generating collision free and dynamically feasible motions remains a central challenge in robotic manipulation, especially in cluttered and non convex environments. Classical sampling based planners explore high dimensional spaces effectively but do not handle dynamics [1, 2], while trajectory optimization and MPC based methods explicitly incorporate smoothness, dynamics and collision constraints [3, 4]. Despite recent progress in collision free MPC, these solvers remain highly sensitive to initialization and often converge to local minima in complex scenes.
Learning-based trajectory priors have recently emerged as a way to address this limitation. In particular, diffusion models [5] have shown strong ability to capture the multimodal structure of feasible trajectories, where several distinct solutions may exist around obstacles. When combined with trajectory optimization, these priors help guide the solver toward promising regions of the search space [6, 7]. However, to generalize to unseen environments, diffusion models must be conditioned on scene information [8], and designing an effective and compact representation of obstacles remains a key challenge.
@@
From this perspective, we first argue that an object-centric representation of the scene provides a compact and highly informative conditioning signal for trajectory diffusion. We propose a Slot-Attention–based object-centric encoding that extracts a small set of latent obstacle descriptors from estimated object poses [9]. This representation captures the essential collision structure while enabling fast inference and improved generalization compared to alternative conditioning strategies. We then show that the resulting scene-conditioned diffusion model interacts naturally with a constraint-based MPC solver, which rigorously and efficiently projects the diffused trajectory prior onto a dynamically feasible and collision-free trajectory. In contrast to previous approaches that rely on simple gradient-based post-processing of diffusion samples, our diffusion prior is refined through full OCP projection running in real time on a physical robot. In addition, we introduce a systematic dataset-generation procedure that combines sampling-based planning with constrained trajectory optimization to produce high-quality supervision trajectories.
@@ II-A Classical and learning-based motion planning methods
Classical sampling-based planners [10, 11], explore high-dimensional configuration spaces effectively but do not account for system dynamics [1, 12]. Trajectory optimization methods [13, 14, 15] address this limitation by explicitly incorporating smoothness and dynamic constraints. MPC further extends this approach, formulating collision avoidance as either a hard constraint or a penalization in the optimal control problem [16, 3, 4], allowing for online planning of dynamically feasible, collision-free motions. Alternative approaches such as Model Predictive Path Integral control [17, 18] and tree-search methods like Monte Carlo Tree Search [19] introduce stochastic exploration to mitigate local minima, but their computational cost often limits real-time applicability. Despite these advances, classical TO, MPC, and sampling-based extensions remain sensitive to initialization and prone to local minima in complex, non-convex environments.
Learning-based methods provide a complementary solution by leveraging prior experience from demonstrations [20, 21] or previously successful trajectories [22, 23, 24, 25], guiding planners or optimizers toward promising regions of the search space. In particular, providing a high-quality warm-start near a feasible solution can dramatically accelerate convergence and improve final solution quality [26, 27, 28]. Achieving such warm-starts is challenging due to the high degree of multi-modality in robotic trajectory generation, where multiple distinct solutions may exist for the same task.
@@ II-B Diffusion models in robotics
Diffusion models [5] have recently emerged as a powerful tool to capture the multi-modal distributions of feasible trajectories. Early work such as Diffuser [29] iteratively denoises trajectories using a learned model and differentiable guiding functions, while Diffusion Motion Planning [30] adds gradient-based guidance for task-specific objectives, such as collision avoidance or smoothness. Extensions incorporating model-based projections [31] further improve the quality of generated trajectories. Composable Diffusion Models [32] learn distributions over constraint-satisfying trajectories and leverage them to handle novel combinations of constraints. While effective in low-dimensional or structured problems, these approaches generally struggle to generalize to cluttered or unstructured scenes. [33] recently applied diffusion models to approximate globally optimal nonlinear MPC. Their approach generates a diverse offline dataset of optimal control sequences, trains a diffusion model to capture the multi-modal distribution of solutions, and samples near-globally optimal trajectories at runtime without relying on initial guesses. While effective in simplified simulated environments, it has not been demonstrated on real robots.
@@ II-C Scene-conditioned diffusion approaches
To improve generalization to unseen environments, several works condition diffusion models directly on scene information [34, 35, 7, 6]. Diffusion Seeder [8] goes a step further by deploying diffusion-generated trajectories on a real robot; however, like prior work such as [30], it generates only open-loop trajectories that are tracked by a low-level controller. These approaches do not perform closed-loop replanning, enforce dynamic feasibility during generation, or integrate with a receding horizon MPC framework. While they improve robustness compared to purely simulated methods, they remain limited to simplified or low-dimensional setups and do not demonstrate real-time, collision-aware control with compliance.
@@ II-D Environment representation
A key factor in scene-conditioned planning is how the environment is represented. Image-based representations [36] can lose geometric fidelity, whereas point clouds [34, 35] capture 3D geometry but are high-dimensional, unstructured, and require specialized networks to extract object-level features. Similarly, Diffusion Seeder [8] relies on depth images, which provide direct 3D information but also suffer from high dimensionality and noise, limiting inference speed and generalization. Structured scene representations, such as PRESTO [6], encode key robot configurations labeled for collision status, while object-centric approaches like Slot-Attention [9] map inputs into distinct object slots. These structured representations yield compact latent spaces, better generalization, and faster inference, especially in cluttered or unseen environments [37]. In this work, we adopt a Slot-Attention-based approach and systematically compare it with image-based, PRESTO, and occupancy grid representations (Sec. IV-C). Building on these principles, our approach integrates object-centric, scene-conditioned diffusion priors with a full-dynamics MPC framework. Unlike prior work, we demonstrate real-time, closed-loop control of a physical robot in cluttered, dynamic environments while strictly enforcing collision and dynamic constraints, highlighting the practical benefits of combining generative priors with model-based control.
@@ III DIFFUSION-GUIDED MPC
Figure 3: Object-centric scene representation using Slot Attention. Given an RGB image of the scene, a pre-trained Slot-Attention encoder [9] extracts a set of object-centric latent embeddings, called slots, each corresponding to a distinct object or obstacle in the environment. These slots are then used to construct a structured representation of the scene, which serves as input to the diffusion-based trajectory generator. This conditioning enables the model to generate trajectories that account for obstacle geometry and layout, while maintaining generalization to unseen environments.
@@ III-A1 Trajectory diffusion
To generate high-quality warm-start trajectories, we train a conditional diffusion model as a learned prior over feasible motion sequences [30, 6]. Our model builds on the Denoising Diffusion Probabilistic Model (DDPM) framework [5], which captures complex, multi-modal distributions through a forward noising and reverse denoising process. A robot trajectory $\bm{Q}\!=\!(\bm{q}_{0},\dots,\bm{q}_{T})$ is treated as a sample from an expert distribution. A forward (noising) process gradually adds Gaussian noise to the trajectory over $K$ steps according to a fixed variance schedule $\{\beta_{i}\}_{i=1}^{K}$. The noisy trajectory at step $i$ is given by:
@@ III-A2 Model Architecture
We use a Diffusion Transformer (DiT) that denoises a trajectory $\bm{\hat{Q}}_{i}$ conditioned on the task context. The context includes the robot’s initial configuration $\bm{q}_{0}$, the Cartesian goal $\bm{p_{\text{goal}}}$, and a scene encoding. The latter is provided by a Slot Attention encoder [9], which converts an RGB image of the workspace into object-centric latent embeddings (slots). This representation captures obstacle information in a compact form, enabling the DiT to generate trajectory proposals that are tailored to each problem instance. We describe the encoder in more detail below.
@@ III-A3 Slot Attention-Based Scene Encoder
The Slot Attention encoder [9] extracts object-centric latent representations from single-view RGB images of the robot workspace. This structured decomposition of the scene into obstacles provides a compact conditioning signal for the diffusion model, enabling generalization across cluttered layouts without relying on fixed spatial encodings or manual annotation.
@@ III-B2 MPC costs and constraints
[…]
d. Collision constraints. To guarantee safety, we impose a hard collision-avoidance constraint at each timestep following [40], requiring the signed-distance $d(\cdot,\cdot)$ between any pair of potentially colliding bodies $\mathcal{B}$ to be greater than a safety margin $\epsilon_{\text{safe}}$:
@@ III-B3 Warm-starting the solver
The performance of the generative model is critically dependent on the quality of its training data. To learn a useful motion prior, the model must be trained on a diverse dataset of expert trajectories that are not only collision-free but also dynamically smooth and efficient. Trajectories from purely geometric planners like RRT [1] are often jerky and not directly executable. Therefore, a dedicated data-generation pipeline is required to produce high-quality examples. We construct our training dataset by generating thousands of random scenes in simulated environments described in Sec. IV. For each scene, we have computed multiple expert trajectories. We first sample a start/goal pair and generate a geometrically valid path using RRT. This path is then used to initialize and solve the OCP defined in (3), resulting in a smooth, feasible expert trajectory. In parallel, the Slot Attention algorithm was trained beforehand on a task-specific dataset. The diffusion model is then trained on this dataset by minimizing the loss in (2).
@@ IV-A2 Metrics and protocol
We evaluate our approach against a range of classical and learning-based motion-planning methods to demonstrate that combining diffusion-based warm-starting with OCP refinement yields superior computational efficiency and higher-quality trajectories. The first learning-based baseline is PRESTO [6], a diffusion model operating in configuration space, where the scene is represented by sampling a predefined set of robot configurations. For each configuration, a binary collision indicator is computed, and the resulting indicator vector is concatenated to form the conditioning input. The second baseline is Motion-Planning Diffusion (MPD) [30], a learning-based approach where an unconditional diffusion model learns to generate motion trajectories, with collision-avoidance encouraged through a cost gradient-based guidance within the diffusion process. For fairness, we adapt both of their architecture and training procedure to match our dataset and task setup. We then compare our approach to both classical and optimization-based planners. RRT Connect [38] is a widely-used bidirectional sampling-based planner, implemented in Python with C++ bindings, that incrementally builds trees from the start and goal configurations and attempts to connect them. CuRobo [17] is a recent GPU-accelerated framework that combines trajectory optimization with sampling strategies to efficiently handle high-dimensional planning problems. To ensure a fair comparison under the same time budget, we terminate all algorithms after a fixed number of iterations.
Finally, to demonstrate the benefit of warm-starting the OCP, we compare our method to a standard OCP solver [39], which solves the full optimal control problem defined in (3) using Differential Dynamic Programming (DDP) [41, 42] without any warm-start. This method directly optimizes dynamic feasibility and task objectives but remains highly sensitive to local minima.
@@
Figure 6: Impact of the cost gradient-based guidance. Comparing the impact of the cost gradient-based guidance at each diffusion step, as used in [30]. Results show that guidance is not helpful with a low computational budget, given by control requirements.
(a) Success rate
(b) Computation time (ms)
@@ IV-C2 Impact of the cost gradient-based guidance
Fig. 6 evaluates the impact of adding a cost gradient-based guidance toward the feasible manifold at each denoising step, following [30]. Without OCP refinement, the cost gradient-based guidance improves success at low budgets by steering trajectories closer to feasibility early on. However, when OCP refinement is applied, its benefit diminishes and even harms performance at higher budgets due to unnecessary overhead. We thus omit the cost gradient-based guidance in our final pipeline.
@@ IV-C3 Trade-off between computation and warm-start quality
Experiments were conducted on a Franka Emika Panda robotic arm controlled via torque commands. The MPC framework was implemented in Python with C++ bindings and executed on an AMD Ryzen 9 5950X CPU at $3.4\text{\,}\mathrm{GHz}$. The dynamics of the system were computed using pinocchio [44], and in our setup the dynamics are time-invariant. The optimal control problem (OCP) was solved using mim-solvers [45] inside crocoddyl [39], consistent with the simulation setup described in Section IV-A. The robot dynamics were modeled using inertial parameters from [46]. Obstacle detection and pose estimation were performed in real-time using AprilTags [47], with known obstacle dimensions and external-camera-based object tracking. Control commands were computed at $1\text{\,}\mathrm{kHz}$, while the OCP was solved at $100\text{\,}\mathrm{Hz}$ with a prediction horizon of $N_{h}=15$ nodes and a time step of ${\Delta t}=$10\text{\,}\mathrm{ms}$\$, resulting in an optimization horizon of $150\text{\,}\mathrm{ms}$. At each OCP iteration, we compute the optimal trajectory and the corresponding Riccati gains [48]. The Riccati gains are used to apply a low-level feedback policy at $1\text{\,}\mathrm{kHz}$ until the next replanning time. A collision safety margin of $1\text{\,}\mathrm{cm}$ was enforced via a lower-bound constraint on the signed-distance. The MPC operated under a receding-horizon scheme. Whenever a new high-level task was issued (e.g., reaching a new target position), the diffusion model generated a warm-start trajectory, replacing the previous solution. As Fig. 8 shows, this warm-started OCP enabled reactive and efficient collision-avoidance in cluttered and dynamically changing environments.